OpenStudy (anonymous):

A 2.50 kg steel ball strikes a wall with a speed of 10.0 m/s at an angle of 60.0° with the surface. It bounces off with the same speed and angle. If the ball is in contact with the wall for 0.230 s, what is the average force exerted on the ball by the wall?

OpenStudy (anonymous):

Average force is the average time rate of change of momentum. since the motion doesn't change in one of the directions (call it the y direction) the average force would be \[F = m \Delta v_{x}/\Delta t = m(v_{fx} - v_{ix})/\Delta t \] \[= 2.5kg(-10m/s \sin(60) - 10 m/s \sin(60))/0.230s \] = 188 N in the negative x direction